# TWIST: Training-free and Label-free Short Text Clustering through Iterative Vector Updating with LLMs

### v1

@@ 5.2 Prompts used
The prompts we used are shown in Table 2. It consists of three parts: task instruction, answer format, and task.
@@
5.3 Hyperparameter setting
The main hyperparameter in our approach is the initial dimension (number of representatives) of the sparse vector $d$, which we set to 2,048 based on validation results on two external datasets: ArxivS2S and Reddit (Muennighoff et al., 2023). We use external datasets to prevent overfitting on a specific dataset. See details in Appendix B. Qwen3-8B (as LLM) and All-MiniLM-L6-v2 (as embedder) are used in the hyperparameter search.
5.4 Evaluation
The emerging and GCD scenarios are evaluated under two conditions: with HDBSCAN (McInnes et al., 2017) and K-means (Lloyd, 1982).
@@ 5.4 Evaluation
To reflect real-world settings with unknown $K$, we apply HDBSCAN (McInnes et al., 2017) to partition $D^{test}$ into $\hat{K}$ clusters. We set min_samples = 1 to reduce noise. The noise points are assigned to the nearest cluster centroid. We perform an incremental grid search over min_cluster_size values of $10,15,\ldots,50$, stopping early if the silhouette score does not improve (Rousseeuw, 1987; Gung et al., 2023).
K-means clustering is the dominant evaluation practice because many studies assume a known cluster number $K$ and incorporate it in their method pipelines Zhang et al. (2023); De Raedt et al. (2023); Liang et al. (2024b); Feng et al. (2024); Zou et al. (2025b). Following their convention, the number of clusters for K-means is set to match the number of ground-truth cluster number $K$.
To reduce cost, we follow (Zhang et al., 2023) and perform a single pipeline run per dataset using a fixed seed to obtain a sparse vector for each text. Most steps are deterministic. The main source of variability comes from splitting candidate examples across multiple LLM calls (e.g., three calls with 10 examples instead of one with 30), done only to save time and GPU memory. LLM outputs are generated via greedy decoding, without sampling. For evaluation, K-means is run over 5 different seeds and averaged. HDBSCAN, which is near-deterministic, is run once. After we derive the clusters, we apply standard clustering metrics for evaluation. These metrics include normalized mutual information (NMI) and clustering accuracy (Acc) (Rand, 1971; Meilă, 2007; Huang et al., 2014; Gung et al., 2023).
5.5 Baselines
Emerging Data Scenario
We compare against SCCL (Zhang et al., 2021), ClusterLLM, (Zhang et al., 2023), IDAS De Raedt et al. (2023), LLMedgeRefine (Feng et al., 2024), and SPILL (Lin et al., 2025). Table 3 shows that our method makes the fewest assumptions and has the lowest requirements. SPILL is the only method under the same assumptions. However, we still include ClusterLLM, IDAS, and LLMedgeRefine, which uses additional information – the known $K$ – in its development pipeline, to provide a more comprehensive comparison.
GCD Scenario
We compare against IDAS De Raedt et al. (2023), Loop (An et al., 2024), ALUP Liang et al. (2024b), IntentGPT Rodriguez et al. (2024) and Glean Zou et al. (2025b). Note that IntentGPT and IDAS can also be implemented only using unlabeled data. Table 3 shows that our method again makes the fewest assumptions and has the lowest requirements.
Emerging data FT known $K$ Labeled data
TWIST (Ours) -
SCCL (Zhang et al., 2021) $\checkmark$ -
ClusterLLM (Zhang et al., 2023) $\checkmark$ $\checkmark$ -
LLMEdgeRefine (Feng et al., 2024) $\checkmark$ -
IDAS (De Raedt et al., 2023) $\checkmark$ -
SPILL (Lin et al., 2025) -
GCD
TWIST (Ours)
Loop (An et al., 2024) $\checkmark$ $\checkmark$ $\checkmark$
ALUP (Liang et al., 2024b) $\checkmark$ $\checkmark$ $\checkmark$
IntentGPT (Rodriguez et al., 2024)
Glean (Zou et al., 2025b) $\checkmark$ $\checkmark$ $\checkmark$
Table 3: Baselines settings. FT: Finetuning. known $K$: Assuming known $K$ during the method pipeline development. Labeled data: Labeled data is required.
6 Results
6.1 Main results
Emerging data scenario
Table 4 shows the HDBSCAN results. Our approach substantially improves performance compared to directly using the embedder. [^1] The estimated number of clusters, $\hat{K}$, is provided in Appendix D. Moreover, the results from BERT (non-contrastive) also improve with our approach. For the K-means results, Table 5 shows that our approach generally outperforms IDAS and ClusterLLM on both metrics across different embedders, and achieves comparable results to LLMEdgeRefine. This indicates that, even without assuming prior knowledge of $K$ during method development – as the compared methods do – our approach remains effective and better reflects real-world practice. While SPILL achieves slightly higher NMI, our method attains substantially higher ACC.
Lastly, we found that the HDBSCAN results overall are similar to those of K-means, indicating that our method remains robust even without knowing $K$.
@@ 6.1 Main results
TWIST (Ours) 84.86 72.24 93.41 86.89 72.53 44.32 75.85 65.12 27.77 29.03 70.88 59.52
Instructor
Plain 82.50 67.34 92.29 79.27 67.63 54.31 67.99 53.54 0.25 11.48 62.13 53.19
TWIST (Ours) 85.63 73.77 93.64 84.62 74.06 46.81 76.52 64.07 31.64 33.65 72.30 60.58
Table 4: Five-benchmark results (Emerging Data Scenario, HDBSCAN). Gemma is used in TWIST. Plain: direct embedding clustering. Bold numbers: best within the embedder.
Bank 77 Clinc150 Mtop Massive GoEmo Average
NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc
Bert
TWIST (Ours) 76.23 (0.14) 60.49 (0.72) 93.25 (0.07) 86.43 (0.65) 71.79 (0.20) 36.41 (0.20) 72.00 (0.13) 60.56 (0.61) 18.66 (0.10) 18.76 (0.29) 66.38 52.53
MiniLM
TWIST (Ours) 85.57 (0.06) 74.07 (0.71) 93.65 (0.07) 86.35 (0.34) 71.92 (0.22) 37.83 (0.72) 76.18 (0.26) 66.15 (0.58) 20.69 (0.09) 24.44 (0.37) 69.60 57.77
E5
SCCL 77.34 (0.62) 63.60 (1.37) 91.89 (0.49) 77.96 (1.78) 70.42 ((0.34) 33.82 ((1.07) 71.57 (0.89) 54.48 (1.80) 20.05 (0.63) 22.03 (0.69) 66.25 50.38
TWIST (Ours) 85.09 (0.19) 74.23 (0.88) 93.31 (0.04) 85.30 (0.58) 71.57 (0.18) 37.32 (1.67) 75.72 (0.22) 62.93 (1.28) 26.58 (0.23) 33.20 (0.54) 70.45 58.60
Instructor
Plain 82.42 (0.68) 65.49 (2.41) 93.27 (0.32) 81.68 (1.31) 71.63 (0.39) 33.14 (0.52) 74.02 (0.54) 54.62 (2.06) 29.18 (0.76) 32.51 (1.51) 70.10 53.49
SCCL 81.77 (1.36) 65.48 (1.36) 92.94 (0.44) 80.85 (0.74) 73.52 (0.38) 34.28 (0.58) 73.90 ((0.36) 54.10 ((1.05) 30.54 (0.68) 34.33 (0.86) 70.53 53.81
ClusterLLM 85.15 (0.41) 71.20 (1.59) 94.00 (0.21) 83.80 (0.41) 73.83 (0.79) 35.04 (0.97) 77.64 (0.21) 60.69 (0.96) 23.89 (0.68) 26.75 (1.76) 70.90 55.50
LLMEdgeRefine - (0.00) - (0.00) 94.86 (0.00) 86.77 (0.00) 72.92 (0.00) 46.00 (0.00) 76.66 (0.00) 63.42 (0.00) 29.74 (0.00) 34.76 (0.00) - -
IDAS - (0.00) - (0.00) 92.35 (0.00) 81.36 (0.00) 72.31 (0.00) 37.30 (0.00) 75.74 (0.00) 63.01 (0.00) 25.57 (0.00) 30.61 (0.00) - -
SPILL 85.01 (0.29) 71.05 (0.83) 93.77 (0.35) 85.14 (1.04) 72.65 (0.32) 37.11 (1.62) 77.62 (0.46) 62.42 (2.06) 32.78 (0.52) 33.21 (2.33) 72.37 57.79
TWIST (Ours) 85.46 (0.18) 74.66 (1.01) 93.92 (0.09) 85.62 (0.71) 72.34 (0.28) 38.40 (0.81) 76.43 (0.03) 63.17 (0.25) 30.62 (0.24) 37.04 (1.55) 71.75 59.78
Table 5: Five-benchmark results (Emerging Data Scenario, K-means). Averages over 5 runs (±SD). Gemma is used in TWIST. Plain: direct embedding clustering. Bold numbers: best within the embedder. SCCL/ClusterLLM/LLMEdgeRefine cited directly from prior work. SCCL results are cited from ClusterLLM; IDAS results are cited from LLMedgeRefine.
GCD scenario
Table 6 shows the HDBSCAN results for GCD. Our method consistently improves the performance of the existing embedder. Table 7 shows the K-means results for GCD. Our approach is on average the best in the unlabeled setting, without any known categories. It even outperforms baselines that know $10\%$ of the categories, and is comparable to other baselines that know $25\%$ of the categories (or up to $75\%$ in the case of IntentGPT). These strong results, together with the fewer assumptions required by our method, highlight its effectiveness and efficiency. Notably, BANK77 and CLINC150 perform similarly with both clustering algorithms, while Stackoveflow shows a noticeable drop.
@@
Table 6: Three-benchmark results (GCD scenario, HDBSCAN) Gemma is used in TWIST. Bold numbers: best. KCR ratio of known category. All-MiniLM-L6-v2 used as backbone embedder. Plain: direct embedding clustering. StackO: StackOverflow dataset.
Bank 77 Clinc150 StackO Average
KCR Method NMI Acc NMI Acc NMI Acc NMI Acc
- Plain 79.08 61.27 89.46 73.73 72.49 72.68 80.34 69.08
$\mathbf{0\%}$
IDAS 82.84 67.43 93.82 85.48 81.26 83.82 85.97 78.91
IntentGPT 81.42 64.22 94.35 83.20 - - - -
TWIST (Ours) 87.67 77.10 95.21 89.07 81.69 82.56 88.19 82.91
$\mathbf{10\%}$
Loop 79.14 64.97 93.52 84.89 75.98 80.50 82.88 76.79
GLEAN 82.23 67.99 95.21 88.71 79.67 82.40 85.70 79.70
$\mathbf{25\%}$
ALUP 84.06 74.61 94.84 88.40 76.58 82.20 85.16 81.74
Loop 83.37 71.40 94.38 86.58 79.10 82.20 85.62 80.06
GLEAN 85.62 76.98 96.27 91.51 80.90 84.10 87.60 84.20
$\mathbf{75\%}$ IntentGPT 85.94 77.21 96.06 88.76 - - - -
Table 7: Three-benchmark results (GCD scenario, K-means). Averages over 5 runs with Gemma. KCR: known categories ratio. All-MiniLM-L6-v2 used as backbone embedder in TWIST. Loop results are cited from GLEAN. Plain: direct embedding clustering.
6.2 Further analysis
Effect of Different LLMs
Our main results were obtained with gemma-2-9b-it. To investigate whether using a different LLM affects performance, Figure 2 shows the results obtained with all-MiniLM-L6-v2 as the embedder and three different LLMs (see Appendix E for details). Overall, the differences between Qwen3-8B and gemma-2-9b-it are similar but an obvious drop with Llama3.1-8B-Instruct. Because Qwen3-8B achieves comparable results, we use it along with all-MiniLM-L6-v2 for remaining analyses to reduce computational cost.
Figure 2: ACC and NMI across all test datasets (Emerging Data and GCD) for three LLMs. Results for K-means are averaged over 5 runs, HDBSCAN from a single run.
Effect of the Number of Closest Neighbors m and Iterative Updates
We use $m=30$ across all the experiment, but this raises the question of whether different values of $m$ would affect performance. Figure 3(a) shows that using too few neighbors leads to lower results. To investigate the effect of iterative updates, Figure 3(b) shows that in subsequent iterations, the results substantially improve, especially after the first iteration. Since most sparse vectors converge within 5 iterations, there are only minimal changes afterward: We find that most datasets have 90% convergence ratio by the 5th iteration (see Appendix F).
Figure 3: Comparison of (a) varying negihbor $m$ and (b) iterative updates on ACC and NMI across all test datasets (Emerging Data and GCD).
A scalable implementation of our method
Our representation relies on LLM guidance for each text, which can be computationally expensive when the dataset is extremely large. To assess scalability, we additionally explore clustering over the full dataset. The goal is to implement our approach on a subset of the texts and then use a model to generalize to the remaining texts (a form of knowledge distillation), without requiring further LLM computation. Specifically, we train a lightweight MLP ($\sim$10M parameters) to map pre-trained embeddings to sparse vectors. See Appendix G for experiment details. Note that we do not use any labeled data. Table 8 shows that our approach still improves the original embeddings by a large margin, which proves the scalability of our method, offering substantial savings in time and computational resources.
7 Conclusions
We propose TWIST (Short Text Clustering through Iterative Sparse Vector Updating with LLMs), an intuitive, domain-adaptive, label-free, and training-free method for short text clustering. Compared with other SOTA methods, TWIST is built under minimal assumptions and requirements. TWIST achieves better or similar results to state-of-the-art approaches without assuming prior knowledge of clusters or labels, making it more aligned with real-world scenarios. Experiments on diverse datasets and smaller LLMs show that our method is robust in low-resource settings, and also scalable to large datasets. In the future, we plan to expand our approach to clustering in specialized domains and explore ways to incorporate human feedback into our method pipeline.
Bank 77 Clinc150 StackOverf Average
Method NMI Acc NMI Acc NMI Acc NMI Acc
Kmeans
Plain 76.31 59.41 88.99 73.99 68.42 65.93 77.91 66.44
TWIST (Ours) 83.80 73.58 92.78 87.27 72.98 76.90 83.19 79.25
HDBSCAN
Plain 76.15 60.27 84.31 62.84 63.01 61.80 74.49 61.64
TWIST (Ours) 83.39 71.79 92.55 86.04 71.08 69.58 82.34 75.80
Table 8: Clustering results on the entire dataset using knowledge distillation. Averages over 5 runs for K-means. Plain: Directly embedding the test dataset with All-MiniLM-L6-v2.
Acknowledgments
Limitations
LLM capability. Our method relies on the abilities of LLMs. Although we empirically show that small LLMs can achieve good results, this is not guaranteed when the dataset comes from a rare domain.
Embedder capability. Although we empirically show that our approach generally improves performance compared to directly using the embedder for clustering, if the embeddings provide a poor initial result, the improvement might be limited.
Appendix A Comparison of Data Splits in GCD scenario: Our Method vs. Prior Work
Table 9 shows the train–validation–test splits used in Loop (An et al., 2024), ALUP (Liang et al., 2024b) and GLEAN Zou et al. (2025b) for the three benchmark datasets. In their setting, a specified ratio of all categories is randomly selected as known categories, referred to as the known category ratio (KCR). For each known category, 10% of the data is chosen to form the labeled dataset [^2] IntentGPT (Rodriguez et al., 2024) mention they use fewer than 10% for each known category, but they do not specify the exact percentage., while the remaining samples constitute the unlabeled dataset. Together, these two subsets form the training dataset. In our experiments, we do not use the validation set and use only 30% of the training dataset, restricted to the unlabeled setting with no access to labeled data. Evaluation is conducted on the same test split.
They train an embedder using the training split and optimizes it based on the labeled validation set, with the test set used solely for evaluation. In contrast, our method performs test-time embedding construction, so it requires to access test data, similar to IDAS De Raedt et al. (2023) and IntentGPT Rodriguez et al. (2024), using all available text data ($D=D^{test}\cup D^{u}$) to construct embeddings.
# clusters # train # valid # test # total
Bank77 77 9,003 1,000 3,080 13,083
Clinc150 150 18,000 2,250 2,250 22,500
StackO 20 18,000 1,000 1,000 20,500
Appendix B Selection of $d$
Table 10 shows the two datasets that we used for selecting the dimensionality of the sparse vector $d$. Figure 4 shows that $1024$, $2048$, and $\min\{4096,N\}$ give similar performance for both K-means and HDBSCAN. We selected value $d=2048$ and used this value across all datasets and experiments.
[…]
ArxivS2S 93 3,674
Reddit 50 3,217
Figure 4: The average ACC and NMI across two datasets for different $d$. Results for K-means are averaged over 5 runs, while HDBSCAN is shown from a single run. $4096$ here means $\min\{4096,N\}$ because the dataset size can be smaller.
Appendix C Prompts for Instructor embedder
Our method is built on top of existing embedders. For instructor-large, a prompt can be added to generate task-specific embeddings. We follow previous research (Su et al., 2022; Liang and Liao, 2023; Lin et al., 2025) and use mostly the same prompts. The prompts we used for the different datasets are listed in Table 11.
Instructor:
Bank77: Represent the bank purpose for retrieval:
Mtop and Stackoverflow: Represent the sentence for retrieval:
Clinc150 and Massive: Represent the sentence for retrieving the purpose:
GoEmo: Represent an emotion sentence for retrieval:
Table 11: The prompts used in Instructor.
Appendix D Estimated Number of Clusters $\hat{K}$ by HDBSCAN
Table 12 and Table 13 show the estimated number of clusters, $\hat{K}$, for the two scenarios. The results indicate that the refined embeddings produced by our method are generally closer to the ground truth than the pre-trained (Plain) embeddings. An exception is the Bank77 dataset.
@@ Appendix D Estimated Number of Clusters $\hat{K}$ by HDBSCAN
TWIST 70 142 82 57 58
Instructor
Plain 71 144 18 29 2
TWIST 84 148 99 58 57
True $K$ 77 150 102 59 27
Table 12: Emerging scenario: Estimated Number of Clusters $\hat{K}$ by HDBSCAN. gemma-2-9b-it used in TWIST.
@@
Table 13: GCD scenario: Estimated Number of Clusters $\hat{K}$ by HDBSCAN. All-MiniLM-L6-v2 used as backbone embedder. gemma-2-9b-it was used in TWIST.
NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc
Kmeans
Gemma 85.57 74.07 93.65 86.35 71.92 37.83 76.18 66.15 20.69 24.44 87.67 77.10 95.21 89.07 81.69 82.56 76.57 67.20
Qwen 85.13 73.18 93.59 86.96 72.23 39.03 74.55 63.64 20.12 22.80 87.04 75.78 94.80 88.21 80.69 82.32 76.02 66.49
Llama 82.46 69.81 91.03 82.21 70.08 36.21 72.47 60.47 17.60 19.67 84.34 70.91 93.83 86.27 75.69 78.68 73.44 63.03
HCBSCAN
Gemma 85.61 75.29 93.86 86.44 74.52 50.36 76.67 67.51 21.63 23.27 87.43 75.16 95.19 89.07 78.54 71.70 76.68 67.35
Qwen 85.04 73.90 93.56 86.42 74.42 50.11 74.39 62.36 22.10 19.81 86.81 75.29 94.78 88.18 78.07 71.50 76.15 65.95
Llama 82.45 68.70 91.00 78.64 72.80 51.12 73.19 63.84 18.53 16.84 84.16 69.35 93.84 84.18 74.78 72.20 73.84 63.11
Table 14: Further analysis of results using different LLMs.
Appendix E Results by Different LLMs
We report the detailed results of all datasets by different LLMs in Table 14. Qwen and Gemma consistently outperform LLaMA. In addition, the average differences between K-means and HDBSCAN are minimal across both metrics for each LLM.
Appendix F Percentage of texts converged at iteration $t$
Appendix G Details of the full dataset clustering
@@
Table 15 shows the number of examples we use with the LLM. The number of examples with LLM guidance is the same as in Table 1, i.e. $|D^{test}|+|D^{u}|$, so we do not need to re-run all experiments. The examples account for 29.6% to 44.2% of datasets. Examples with LLM guidance are used to train the MLP, since their sparse vectors have been derived. The pre-trained embeddings from all-MiniLM-L6-v2 serve as input, and the LLM-generated sparse vectors serve as target output for supervised learning with an MLP. 20% of the examples are used as a validation set to tune the hyperparameters.
@@
We first encode all examples into pre-trained embeddings using all-MiniLM-L6-v2, then use the trained MLP to derive sparse versions of these embeddings. These sparse vectors are used to cluster all examples.

### v2

@@ Appendix E Prompt
Table 8 shows the prompts used in our experiment. It consists of three parts: task instruction, answer format, and task.
@@
Appendix F Effect of Different Embedders
Because we use a subset instead of scanning the entire dataset for each text, the quality of the texts selected by the embedder influences the construction of our BoT vector. We report two ratios of the first iteration, averaged across all texts.
Embedder Selection Ratio:
$\frac{\#\{\text{True same-cluster in top $m$ as the text}\}}{m}$
In our experiments, we set $m=30$. Note that Embedder here does not refer to the pre-training embeddings alone; it denotes the concatenation of the pre-training embeddings and the BoT vectors. We focus on the first iteration, i.e., $\mathbf{U}^{1}$, since all BoT vectors have not yet converged and they are largely evenly spaced and uninformative in the initial state. As a result, the selection relies mostly on the pre-training embeddings.
LLM Selection Ratio (LLM):
[…]
Table 9 shows that E5 produces the best pool on average. We also observe that the LLM ratio remains relatively stable, even when the embedder ratio fluctuates substantially.
Bank 77 Clinc150 Mtop Massive GoEmo Average
Emb LLM Emb LLM Emb LLM Emb LLM Emb LLM Emb LLM
TF-IDF 43.26 68.02 49.71 82.60 62.93 78.55 45.28 70.21 30.85 36.32 46.41 67.14
Bert 38.08 63.41 56.12 86.49 69.72 82.48 50.48 73.61 20.02 25.90 46.88 66.38
MiniLM 62.03 75.55 69.76 89.09 70.88 81.27 61.58 76.14 22.33 27.79 57.32 69.97
E5 58.64 73.25 67.80 88.48 72.00 81.53 61.66 76.33 27.97 32.02 57.61 70.32
Table 9: Selection Accuracy By Embedders at 1st iteration ($\%$)(emerging data scenario). Bold numbers: best. LLM is Gemma.
Appendix G Percentage of Texts Converged at Iteration $t$
NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc
Gemma 85.44 73.12 94.45 87.58 72.18 39.40 76.74 66.84 21.77 25.22 86.90 74.93 95.32 88.71 81.21 82.54 76.75 67.29
Qwen 85.01 71.63 93.79 87.12 73.29 37.45 74.24 61.47 21.61 25.83 86.53 74.31 95.61 89.41 82.15 83.42 76.53 66.33
Llama 81.75 66.18 91.66 83.12 70.12 35.79 73.26 60.40 18.91 20.81 83.16 68.16 94.00 85.66 77.42 80.10 73.78 62.53
Table 10: Further analysis of results using different LLMs.
Appendix H Results by Different LLMs
We report the detailed results of all datasets by different LLMs in Table 10. Qwen and Gemma consistently outperform LLaMA. We also report the LLM selection accuracy (see Appendix F for the definition). Table 11 shows that Gemma and Qwen have similar capabilities in identifying similar texts, outperforming LLaMA. This finding aligns with prior work Lin et al. (2025).
Bank 77 Clinc150 Mtop Massive GoEmo Average
Emb LLM Emb LLM Emb LLM Emb LLM Emb LLM Emb LLM
Gemma 62.03 75.55 69.76 89.09 70.88 81.27 61.58 76.14 22.33 27.79 57.32 69.97
Qwen 61.08 75.32 70.16 90.14 72.44 84.19 60.88 76.62 22.27 28.91 57.36 71.04
Llama 56.15 64.57 64.05 78.95 67.99 75.16 58.17 67.74 19.96 23.43 53.26 61.97
Table 11: Selection Accuracy By LLMs at 1st iteration($\%$)(emerging data scenario). Bold numbers: best. All-MiniLM-L6-v2 used as backbone embedder.
Appendix I HDBSCAN Clustering
@@
To reflect real-world settings with unknown $K$, we apply HDBSCAN (McInnes et al., 2017) to partition $D^{test}$ into $\hat{K}$ clusters. We set min_samples = 1 to reduce noise. The noise points are assigned to the nearest cluster centroid. We perform an incremental grid search over min_cluster_size values of $10,15,\ldots,50$, stopping early if the silhouette score does not improve (Rousseeuw, 1987). Using the silhouette score for hyperparameter search follows prior work (Gung et al., 2023).
Tables 12 and 13 shows the detailed evaluation result with HDBSCAN clustering.
@@
LeBoT (Ours) 85.03 70.65 93.82 85.80 75.38 53.17 76.33 62.59 28.07 29.59 71.73 60.36
Table 12: Five-benchmark results (emerging data scenario, HDBSCAN). Gemma is used in LeBoT. Plain: direct embedding clustering. Bold numbers: best within the embedder.
@@
Table 13: Three-benchmark results (GCD scenario, HDBSCAN) Gemma is used in LeBoT. Bold numbers: best. All-MiniLM-L6-v2 used as backbone embedder. Plain: direct embedding clustering.
Table 14 and Table 15 show the estimated number of clusters, $\hat{K}$, for the two scenarios. The results indicate that the refined embeddings produced by our method are generally closer to the ground truth than the pre-trained (Plain) embeddings. Except for the Bank77 and GoEmo datasets.
@@
LeBoT 68 147 64 63 60
Table 14: Emerging scenario: Estimated Number of Clusters $\hat{K}$ by HDBSCAN. gemma-2-9b-it used in LeBoT.
@@
Table 15: GCD scenario: Estimated Number of Clusters $\hat{K}$ by HDBSCAN. All-MiniLM-L6-v2 used as backbone embedder. gemma-2-9b-it was used in LeBoT.
Appendix J Details of the full dataset clustering
@@
Table 16 shows the number of examples we use with the LLM. We re-use the representations derived in the emerging data setting, so the number of examples with LLM guidance remains the same as in Table 1, i.e., $|D^{test}|$, and there is no need to re-run all experiments. The examples account for 20.6% to 23.5% of datasets. Examples with LLM guidance are used to train the MLP, since their BoT vectors have been derived. The pre-trained embeddings from all-MiniLM-L6-v2 serve as input, and the LLM-generated BoT vectors serve as target output for supervised learning with an MLP. 20% of the examples are used as a validation set to tune the hyperparameters.
@@
We first encode all examples into pre-trained embeddings using all-MiniLM-L6-v2, then use the trained MLP to derive BoT versions of these embeddings. These BoT vectors are used to cluster all examples.